# Feedback Friction: LLMs Struggle to Accurately Incorporate External Feedback

### v1

@@ Abstract
Recent studies have shown LLMs possess some ability to improve their responses when given external feedback. However, it remains unclear how effectively and thoroughly these models can incorporate extrinsic feedback. In an ideal scenario, if LLMs receive near-perfect and complete feedback, we would expect them to fully integrate the feedback and change their incorrect answers to correct ones. In this paper, we systematically investigate LLMs’ ability to incorporate feedback by designing a controlled experimental environment. For each problem, a solver model attempts a solution, then a feedback generator with access to near-complete ground-truth answers produces targeted feedback, after which the solver tries again. We evaluate this pipeline across a diverse range of tasks, including math reasoning, knowledge reasoning, scientific reasoning, and general multi-domain evaluations with state-of-the-art language models including Claude 3.7 (with and without extended thinking). Surprisingly, even under these near-ideal conditions, solver models consistently show resistance to feedback, a limitation that we term Feedback Friction. To mitigate this limitation, we experiment with sampling-based strategies like progressive temperature increases and explicit rejection of previously attempted incorrect answers, which yield improvements but still fail to help models achieve target performance. We also perform a rigorous exploration of potential causes of Feedback Friction, ruling out factors such as model overconfidence and data familiarity. We hope that highlighting this issue in LLMs and ruling out several apparent causes will help future research in self-improvement.
Figure 1: Top: Accuracy of various solver models when iteratively exposed to feedback of a feedback model (GPT-4.1 mini) with access to ground-truth answers. The horizontal dotted line represents the target accuracy models could theoretically attain if they successfully incorporated all feedback (details in §4.1). Despite receiving high-quality feedback, solver models consistently plateau below their target accuracy. Bottom: Breakdown of questions that remained unsolved by the strongest solver model tested (Claude 3.7 Thinking) after multiple correction attempts. Feedback resistance, rather than feedback quality issues, is responsible for the majority of persistent errors.
@@ 1 Introduction
The prospect of self-improving Large Language Models (LLMs) has sparked both excitement and debate. Questions persist about LLMs’ inherent ability to self-improve without external guidance [1, 2], yet several studies have shown that LLMs can boost their performance when provided with accurate external feedback during test time without any parameter updates or training [3, 4, 5]. However, while prior studies establish the existence of performance gains from external feedback, the upper bounds of such improvement remains largely unexplored. This improvement capability would have far-reaching implications for applications such as scientific discovery [6, 7, 8] and complex planning [9, 10], where repeated refinement cycles with feedback are critical for success. This raises a critical question: can models fully incorporate correct feedback to self-improve and reach their maximum potential?
@@
Our analysis in §4 reveals several key insights into Feedback Friction. First, we categorize errors that persist after multiple feedback iterations and find that feedback resistance (i.e., models failing to incorporate clear and accurate feedback) is the dominant failure mode across all tasks4.1). Second, we attempt to mitigate it by avoiding repeated wrong answers through sampling strategies. While performance improves across tasks, models still fall below their target accuracy (§4.2). Finally, we investigate why models resist feedback (§4.3). Our investigations rule out several apparent causes, including model confidence, reasoning complexity, and data familiarity (specifically, whether frequently occurring entities in pre-training data affect receptiveness to feedback). We also examine whether the same questions consistently challenge models, but surprisingly, different models struggle with different subsets. This suggests that resistance to feedback is not solely determined by the target datasets. In summary, having shown that state-of-the-art LLMs consistently resist external feedback, we then explore and reject the most apparent explanations for this phenomenon.  [^1] Code for this work is available at: https://github.com/JHU-CLSP/Feedback-Friction
@@ 3.1 Experimental setup
We employ nine diverse tasks for evaluation, deliberately choosing objective tasks with clear ground-truth answers to ensure reliable evaluation of feedback incorporation, as using another LLM to evaluate more subjective tasks like instruction following or translation could lead to issues like reward hacking and unreliable assessments. Our tasks include: AIME 2024 [11] and MATH-500 [12] for mathematical problem-solving, TriviaQA [13] and PopQA [14] for knowledge reasoning, MMLU [15] and MMLU Pro [16] for multi-domain evaluation, GPQA [17] for complex scientific reasoning, and two synthetic digit multiplication tasks. The first synthetic task involves 5-digit multiplication (e.g., 78934 × 62851), while the second task applies hexadecimal multiplication rules to decimal numbers (i.e., first mapping 0-9 to themselves and 10-15 to their hexadecimal digits ($10\rightarrow A$, …, $15\rightarrow F$), then performing the arithmetic as if operating in base-16—creating a counterfactual [18] arithmetic setting that challenges models’ learned numerical reasoning patterns. For both tasks, ground-truth solutions are generated using deterministic templates that break down the multiplication into smaller multiplication and addition operations. We include these synthetic tasks specifically to remove potential confounding variables from semantic context (more details about these tasks can be found in Appendix C).
@@
Other than the two synthetic tasks, only AIME 2024, GPQA, and MATH-500 include complete solutions, while the others provide only final answers. For MMLU and MMLU Pro, the multiple-choice format raises concerns about whether models might solve problems through simple elimination strategies rather than genuine reasoning (e.g., selecting A in the first iteration, B in the second, C in the third, etc). However, our analysis reveals that models exhibit surprising choice persistence even when provided with corrective feedback. Rather than systematically eliminating options or switching between choices, models often remain anchored to their initial selections (typically one or two specific answer choices) across multiple feedback iterations, even when that choice is demonstrably incorrect. We use the same prompts, few-shot demonstrations, answer parsing mechanism, and metrics from lm-evaluation-harness and llama-evaluate where applicable. For MMLU, MMLU Pro, PopQA, and TriviaQA, we sample 10% of the total data points, as running the full dataset through our 10-iteration improvement process would be prohibitively time-consuming. Our preliminary experiments confirmed that this subset yields performance metrics nearly identical to those obtained from the complete dataset.
Preventing answer leakages in feedback messages
@@
Beyond the feedback mechanisms discussed in §2.2, we implement safeguards to prevent information leakage by ensuring feedback does not directly reveal the final correct answer. Specifically, we add a filter mechanism to mask out the final answer in the feedback only when the complete correct answer is detected through regex pattern matching. This targeted masking preserves the intelligibility of feedback while preventing direct answer revelation. For example, in mathematical tasks, we replace standalone numerical answers with “[masked]” (e.g., “The final answer is [masked]” instead of “The final answer is 42”) while preserving intermediate steps and reasoning guidance  [^2] We verified that models do not attempt to predict the “[masked]” token during inference. This approach ensures that feedback remains informative and comprehensible while maintaining evaluation integrity. Detailed prompts for problem solving, feedback generation, and concrete examples of the masking process can be found in Appendix A.
@@
For the Llama models, during inference, we use temperature 0 to ensure deterministic outputs and conduct inference using vLLM [22] with each model’s corresponding chat template. All inference is performed on a single H100 instance equipped with eight 80GB GPUs. For Claude models, we access them through Anthropic’s API. For Claude 3.7, we use temperature 0 to ensure deterministic outputs, while for Claude 3.7 with extended thinking, we use temperature 1 as suggested by Claude.
@@ 3.2 Main findings
Figure 3 shows results using our strongest feedback mechanism (Strong-Model Reflective Feedback ($F_{3}$)) across all datasets. Our results reveal a striking pattern: Despite receiving high-quality feedback, all solver models consistently plateau below their target accuracy, which is the accuracy of an ideal model if it successfully incorporates all the given feedback (see §4.1 for calculation details). While Claude 3.7 Thinking achieves the highest initial accuracy on several tasks (particularly AIME, TriviaQA, GPQA, and MATH-500), and Claude 3.7 shows competitive performance across most benchmarks, both Claude variants exhibit the same fundamental plateauing behavior as the Llama models. The performance typically improves rapidly through the first 2-4 iterations before significantly slowing down. This Feedback Friction is particularly pronounced on complex reasoning tasks like AIME and GPQA, where even the best-performing models remain 15-25% and 3-8% below their respective theoretical ceilings despite 10 correction opportunities. Interestingly, Claude 3.7 Thinking shows strong initial performance on AIME (starting around 50% vs. 20-35% for other models) but still plateaus well below the target, suggesting that even sophisticated reasoning processes struggle with iterative feedback incorporation. The synthetic tasks reveal particularly interesting patterns across model families. In the standard 5-Digit Multiplication task, both Claude models reach near-perfect accuracy after significant initial improvement, outperforming the Llama models. However, the Hexadecimal-5-Digit-Multiplication task reveals extreme difficulty with feedback incorporation across all models—no model exceeds 20% accuracy even after 10 iterations, highlighting severe limitations in feedback integration for counterfactual arithmetic systems.  [^3] While it is true that models’ imperfect understanding of hexadecimal arithmetic leads to imperfect feedback quality, this is reflected in the lower theoretical ceiling shown in the figure. Even accounting for this limitation, models still fall short of what they could achieve if they fully incorporated the available feedback.
@@
Feedback quality significantly impacts performance gains. Figure 4 illustrates how different feedback mechanisms affect performance across models and tasks. Due to the cost of running extensive experiments with Claude models, we focus the feedback quality analysis on the Llama model families. All tasks show clear benefits from increasingly sophisticated feedback, with a consistent progression at the final iteration (K=10).
@@
The impact of high-quality feedback is most pronounced on complex reasoning tasks. For AIME, MMLU Pro, and GPQA, Strong-Model Reflective Feedback ($F_{3}$) outperforms binary feedback by significant margins across all models. Llama-4-Maverick shows the strongest overall performance, achieving 73.3% accuracy on AIME and 96.5% on GPQA with Strong-Model Reflective Feedback ($F_{3}$)—-improvements of +26.7% and +10.6% over binary feedback, respectively. Llama-4-Scout demonstrates the largest relative gains, particularly on AIME (+33.3%) and GPQA (+13.1%), compared to Llama-3.3 which shows more modest improvements (+26.7% on AIME, +6.6% on GPQA). This variation in improvement rates suggests that newer and more powerful models may be more receptive to high-quality feedback. Nevertheless, despite these substantial improvements, all models still plateau significantly below their theoretical performance ceiling.
@@ 4.1 Feedback integration failures dominate persistent self-improvement errors
Table 1 presents the distribution of error categories across different tasks, as classified by o4-mini. As shown, feedback resistance is consistently the dominant category across all tasks, accounting for 62.8-100% of errors. This finding suggests that the core challenge in achieving perfect performance lies not in the quality of feedback or problem complexity, but in fundamental limitations of how models process and incorporate corrective feedback. Detailed examples of each error category are provided in Appendix B.
@@ 4.2 Mitigating Feedback Friction with sampling strategies
As shown in Figure 5, the combined strategy yields substantive performance gains across both multiple-choice datasets and non-multiple-choice datasets compared to the baseline that only increases the temperature. While the magnitude of these improvements varied, the consistent pattern suggests that forcing models to explore new solution paths by rejecting previously used answers is beneficial regardless of the task format. Despite these gains, we observed that all datasets still fall short of the target accuracy, indicating that sampling strategies alone cannot fully resolve model resistance to feedback.
@@ 4.3 Understanding Feedback Friction
Figure 6: Accuracy of Llama4 Scout (top) and Llama4 Maverick (bottom) on 5-digit multiplication.
@@
Models with high confidence in their answers may resist feedback and correction, so does excessive confidence cause Feedback Friction? To test this, we measure confidence using the average probability of tokens in the model’s output in the first round, following a similar methodology that has been proven effective in information retrieval domains [25]. We conduct this experiment on 5-digit multiplication tasks to eliminate confounding variables from semantic context. The results shown in Figure 6 indicate no significant correlation between confidence metrics and final accuracy on Llama-4-Scout and Llama-4-Maverick. Higher confidence in initial answers neither predicts poor performance after self-improvement iterations nor correlates with the magnitude of improvement. Additional detailed analysis is available in Appendix D, including experiments with more models and across a wider range of datasets that further support this finding.
@@
We use Llama-3.3 for this experiment since it was released closer to the publication of PopQA. These popularity metrics provide a particularly relevant measure of potential training data frequency for this model. As shown in Figure 7, we found no consistent pattern between entity popularity and accuracy. We provide further supporting evidence in Appendix E with additional statistical testing and alternative popularity metrics.
@@
We also explored whether problem complexity, as measured by the number of expected reasoning steps, correlates with Feedback Friction since prior works [27] show that longer reasoning trajectories may impede problem-solving in LLMs. Additionally, we investigated whether certain questions consistently induced “stubbornness” across different models because some questions may inherently be more difficult to answer. For both hypotheses, our controlled experiments yielded negative results, with minimal correlation between reasoning complexity and stubbornness, and little overlap in stubborn questions across models. Complete analyses are provided in Appendix F and Appendix G.

### v2

# Feedback Friction: LLMs Struggle to Fully Incorporate External Feedback

@@ Abstract
Recent studies have shown LLMs possess some ability to improve their responses when given external feedback. However, it remains unclear how effectively and thoroughly these models can incorporate extrinsic feedback. In an ideal scenario, if LLMs receive near-perfect and complete feedback, we would expect them to fully integrate the feedback and reach correct solutions. In this paper, we systematically investigate LLMs’ ability to incorporate feedback by designing a controlled experimental environment. For each problem, a solver model attempts a solution, then a feedback generator with access to near-complete ground-truth answers produces targeted feedback, after which the solver tries again. We evaluate this pipeline across a diverse range of tasks, including math reasoning, knowledge reasoning, scientific reasoning, and general multi-domain evaluations with state-of-the-art language models including Claude 3.7 with extended thinking. Surprisingly, even under these near-ideal conditions, solver models consistently show resistance to feedback, a limitation that we term Feedback Friction. To mitigate this limitation, we experiment with sampling-based strategies like progressive temperature increases and explicit rejection of previously attempted incorrect answers, which yield improvements but still fail to help models achieve target performance. We analyze Feedback Friction and find that models’ confidence on specific questions, measured by semantic entropy, predicts feedback resistance: high-confidence predictions remain resistant to external correction. We hope that highlighting this issue in LLMs will help future research in self-improvement.
Figure 1: Top: Accuracy of various solver models when iteratively exposed to feedback from a feedback model (GPT-4.1 mini) with access to ground-truth answers. The horizontal dotted line represents the target accuracy models could theoretically attain if they successfully incorporated all feedback (details in §4.1). Despite receiving high-quality feedback, solver models consistently plateau below their target accuracy. Bottom: Breakdown of questions that remained unsolved by the strongest solver model tested (Claude 3.7 Thinking) after multiple correction attempts. Feedback resistance, rather than feedback quality issues, is responsible for the majority of persistent errors.
@@ 1 Introduction
The prospect of self-improving Large Language Models (LLMs) has sparked both excitement and debate. Questions persist about LLMs’ inherent ability to self-improve without external guidance [1, 2], yet several studies have shown that LLMs can boost their performance when provided with accurate external feedback during test time without any parameter updates or training [3, 4, 5]. However, while prior studies establish the existence of performance gains from external feedback, the upper bounds of such improvement remains largely unexplored. The extent of this improvement would have far-reaching implications for applications such as scientific discovery [6, 7, 8] and complex planning [9, 10], where repeated refinement cycles with feedback are critical for success. This raises a critical question: can models fully incorporate correct feedback to self-improve and reach their maximum potential?
@@
Our analysis in §4 reveals several key insights into Feedback Friction. First, we categorize errors that persist after multiple feedback iterations and find that feedback resistance (i.e., models failing to incorporate clear and accurate feedback) is the dominant failure mode across all tasks4.1). Second, we attempt to mitigate it by avoiding repeated wrong answers through sampling strategies. While performance improves across tasks, models still fall below their target accuracy (§4.2). Finally, we investigate why models resist feedback (§4.3). Our investigations rule out several apparent causes, including reasoning complexity, data familiarity, and whether specific questions consistently challenge all models. However, we uncover two key findings: First, using semantic entropy to measure confidence, we find that less confident models show greater relative improvements from feedback, indicating confident models are more resistant to correction. Second, models consistently claim to understand feedback and express willingness to update their beliefs (>95%) yet fail to actually incorporate corrections—revealing a disconnect between stated intentions and actual behavior. In summary, having shown that state-of-the-art LLMs consistently resist external feedback, we identify important patterns in this resistance while ruling out several intuitive explanations. [^1] Code for this work is available at: https://github.com/JHU-CLSP/Feedback-Friction
@@ 3.1 Experimental setup
We employ nine diverse tasks for evaluation, deliberately choosing objective tasks with clear ground-truth answers to ensure reliable evaluation of feedback incorporation.  [^2] Using another LLM to evaluate more subjective tasks like instruction following or translation could lead to issues like reward hacking and unreliable assessments. Our tasks include: AIME 2024 [11] and MATH-500 [12] for mathematical problem-solving, TriviaQA [13] and PopQA [14] for knowledge reasoning, MMLU [15] and MMLU Pro [16] for multi-domain evaluation, GPQA [17] for complex scientific reasoning, and two synthetic digit multiplication tasks. The first synthetic task involves 5-digit multiplication (e.g., 78934 × 62851), while the second task applies hexadecimal multiplication rules to decimal numbers (i.e., first mapping 0-9 to themselves and 10-15 to their hexadecimal digits ($10\rightarrow A$, …, $15\rightarrow F$), then performing the arithmetic as if operating in base-16—creating a counterfactual [18] arithmetic setting that challenges models’ learned numerical reasoning patterns. For both tasks, ground-truth solutions are generated using deterministic templates that break down the multiplication into smaller multiplication and addition operations. We include these synthetic tasks specifically to remove potential confounding variables from semantic context (more details about these tasks can be found in Appendix C).
@@
Other than the two synthetic tasks, only AIME 2024, GPQA, and MATH-500 include complete solutions, while the others provide only final answers. For MMLU and MMLU Pro, the multiple-choice format raises concerns about whether models might solve problems through simple elimination strategies rather than genuine reasoning (e.g., selecting A in the first iteration, B in the second, C in the third, etc). However, our analysis reveals that models exhibit surprising choice persistence even when provided with corrective feedback. Rather than systematically eliminating options or switching between choices, models often remain anchored to their initial selections (typically one or two specific answer choices) across multiple feedback iterations, even when that choice is demonstrably incorrect. We use the same prompts, few-shot demonstrations, answer parsing mechanism, and metrics from lm-evaluation-harness and llama-evaluate where applicable to maintain established evaluation practices and enable fair comparison with prior work. For MMLU, MMLU Pro, PopQA, and TriviaQA, we sample 10% of the total data points, as running the full dataset through our 10-iteration improvement process would be prohibitively time-consuming. Our preliminary experiments confirmed that this subset yields performance metrics nearly identical to those obtained from the complete dataset.
Prompt design and experimental controls
Our experimental framework employs carefully structured prompts to ensure consistent feedback delivery across iterations. For the solver model, we use task-specific system prompts (detailed in Appendix A) and construct iterative prompts that include complete interaction history. We investigate whether prompt structure and organization affect feedback incorporation by comparing our standard single-prompt approach against a multi-turn conversation format that structures the same information across multiple dialogue exchanges. In the conversation format, we reformatted the iterative improvement process to mimic natural dialogue, with alternating turns between solver attempts (as user) and feedback responses (as assistant). Results showed marginal differences compared to single-prompt formatting, suggesting that Feedback Friction persists regardless of interaction structure.
@@
Beyond the feedback mechanisms discussed in §2.2, to maintain evaluation integrity while preserving feedback quality, we implement comprehensive answer masking to prevent feedback from directly revealing ground truth solutions. We use “[masked]” as the replacement token for filtered content, applying targeted masking that preserves intermediate steps and reasoning guidance while preventing direct answer revelation. For example, we replace numerical answers with “[masked]” (e.g., “The final answer is [masked]” instead of “The final answer is 42”) while preserving solution steps  [^3] We verified that models do not attempt to predict the “[masked]” token during inference.
@@
For the Llama models, during inference, we use temperature 0 to ensure deterministic outputs and conduct inference using vLLM [22] with each model’s corresponding chat template. All inference is performed on a single H100 instance equipped with eight 80GB GPUs. For Claude models, we access them through Anthropic’s API. For Claude 3.7, we use temperature 0 to ensure deterministic outputs, while for Claude 3.7 with extended thinking, we use temperature 1 as suggested by Claude. We also explore the effects of varying these temperature settings in our experiments to mitigate Feedback Friction4.2).
@@ 3.2 Main findings
Figure 3 shows results using our strongest feedback mechanism (Strong-Model Reflective Feedback ($F_{3}$)) across all datasets and all models. Our results reveal a striking pattern: Despite receiving high-quality feedback, all solver models consistently plateau below their target accuracy, which is the accuracy of an ideal model if it successfully incorporates all the given feedback (see §4.1 for calculation details). Claude 3.7 Thinking achieves the highest initial accuracy on several tasks (AIME, TriviaQA, GPQA, and MATH-500), while Claude 3.7 shows competitive performance across most benchmarks. However, both Claude variants exhibit the same fundamental plateauing behavior as the Llama models. The performance typically improves rapidly through the first 2-4 iterations before significantly slowing down. This Feedback Friction is particularly pronounced on complex reasoning tasks like AIME and GPQA, where even the best-performing models remain 15-25% and 3-8% below their respective theoretical ceilings despite 10 correction opportunities. The synthetic tasks reveal particularly interesting patterns across model families. In the standard 5-Digit Multiplication task, both Claude models reach near-perfect accuracy after significant initial improvement, outperforming the Llama models. However, the Hexadecimal-5-Digit-Multiplication task reveals extreme difficulty with feedback incorporation across all models—no model exceeds 20% accuracy even after 10 iterations, highlighting severe limitations in feedback integration for counterfactual arithmetic systems.  [^4] While it is true that models’ imperfect understanding of hexadecimal arithmetic leads to imperfect feedback quality, this is reflected in the lower theoretical ceiling shown in the figure. Even accounting for this limitation, models still fall short of what they could achieve if they fully incorporated the available feedback.
@@
Feedback quality significantly impacts performance gains. Figure 4 illustrates how different feedback mechanisms affect performance across models and tasks. Due to the cost of running extensive experiments with Claude models, we focus the feedback quality analysis on the Llama model families. All tasks show clear benefits from increasingly sophisticated feedback.
@@
The impact of high-quality feedback is most pronounced on complex reasoning tasks. For AIME, MMLU Pro, and GPQA, Strong-Model Reflective Feedback ($F_{3}$) outperforms binary feedback by significant margins across all models. Llama-4-Maverick shows the strongest overall performance, achieving 73.3% accuracy on AIME and 96.5% on GPQA with Strong-Model Reflective Feedback ($F_{3}$)—-improvements of +26.7% and +10.6% over binary feedback, respectively. Llama-4-Scout demonstrates the largest relative gains, particularly on AIME (+33.3%) and GPQA (+13.1%), compared to Llama-3.3 which shows more modest improvements (+26.7% on AIME, +6.6% on GPQA). Nevertheless, despite these substantial improvements, all models still plateau significantly below their theoretical performance ceiling.
@@ 4.1 Feedback integration failures dominate persistent self-improvement errors
Table 1 presents the distribution of error categories across different tasks, as classified by o4-mini. Feedback resistance is consistently the dominant category across all tasks, accounting for 62.8-100% of errors. This finding suggests that the core challenge in achieving perfect performance lies not in the quality of feedback or problem complexity, but in fundamental limitations of how models process and incorporate corrective feedback. Detailed examples of each error category are provided in Appendix B.
@@ 4.2 Mitigating Feedback Friction with sampling strategies
As shown in Figure 5, the combined strategy yields substantive performance gains across both multiple-choice datasets and non-multiple-choice datasets compared to the baseline that only increases the temperature. While the magnitude of these improvements varied, the consistent pattern suggests that forcing models to explore new solution paths by rejecting previously used answers is beneficial. Despite these gains, we observed that all datasets still fall short of the target accuracy, indicating that sampling strategies alone cannot fully resolve model resistance to feedback.
@@ 4.3 Understanding Feedback Friction
Could excessive model confidence explain resistance to feedback in Feedback Friction? To test this, we measure confidence using semantic entropy [25], a method that captures uncertainty at the meaning level rather than surface form variations. Unlike token-level probability measures, semantic entropy accounts for the fact that models can express the same meaning in multiple ways, providing a more robust measure of true uncertainty.
To compute semantic entropy for each model response, we first generate multiple outputs $\{s_{1},s_{2},\ldots,s_{n}\}$ by sampling the model $n=50$ times with temperature $0.7$, where each $s_{i}$ represents a complete answer given the feedback. We then cluster these outputs using final answers, which we find to have similar results as semantic equivalence grouping using bidirectional entailment. For each semantic cluster $C_{i}$, we estimate its probability as $\hat{P}(C_{i}|x)=\sum_{s\in C_{i}}P(s|x)/\sum_{j=1}^{K}\sum_{s\in C_{j}}P(s|x)$, where $x$ represents the input (question, previous answer, and feedback), and compute semantic entropy as $H=-\sum_{i}\hat{P}(C_{i}|x)\log\hat{P}(C_{i}|x)$. Lower entropy indicates the model consistently generates semantically equivalent answers (high confidence), while higher entropy suggests diverse semantic meanings (low confidence).
We conduct this analysis across 5 digits multiplication, GPQA, MATH, MMLU-Pro, and TriviaQA using semantic entropy with bucket size 0.2. Figure 6 shows some of the results (full results in Appendix I) and reveals a consistent pattern: absolute improvement rate (the green trends) increases with semantic entropy, rising from near-zero at low entropy (high confidence) to 0.4-0.8 at higher entropy levels (low confidence). These results suggest that models with lower confidence (higher semantic entropy) have both more room for improvement and greater receptiveness to feedback, while highly confident models show minimal improvement despite receiving the same quality feedback.
Figure 6: Relationship between semantic entropy and feedback incorporation across three benchmark tasks (MATH, MMLU-Pro, GPQA) for Llama-3.3-70B and Llama-4-Scout models. Semantic entropy (x-axis) measures model uncertainty, with lower values indicating higher confidence. Three metrics are shown: initial accuracy before any feedback (blue), final accuracy after iterative feedback (orange), and absolute improvement rate calculated as (FinalInitial)/(1Initial)(Final-Initial)/(1-Initial) (green). Numbers in boxes indicate sample sizes per bucket. The absolute improvement rate generally increases with semantic entropy across all tasks and models, indicating that models with lower initial confidence (higher entropy) show greater relative improvements from feedback.
Model self-perception versus actual behavior. To probe deeper into Feedback Friction mechanisms, we directly asked solver models about their understanding of feedback and willingness to incorporate it. After receiving feedback, we prompt models with: (1) “Do you understand this feedback?” and (2) “Will you update your belief or understanding about this problem?”
Surprisingly, across all models and tasks, solver models consistently claimed to understand the feedback ($95\%$ “yes” responses) and expressed willingness to update their beliefs ($96\%$ “yes” responses). However, in subsequent iterations, these same models failed to actually incorporate the feedback. This disconnect between stated intention and actual behavior reveals a fundamental issue: models exhibit self-assessment failure where they believe they are incorporating feedback while demonstrably failing to do so (the prompts used for probing and example conversations can be found in Appendix H). This disconnect between stated intention and actual behavior reveals a fundamental implementation failure: models express both understanding and willingness to change but fail to execute these changes in practice.
@@
We use Llama-3.3 for this experiment since it was released closer to the publication date of PopQA. These popularity metrics provide a particularly relevant measure of potential training data frequency for this model. As shown in Figure 7, we found no consistent pattern between entity popularity and accuracy. We provide further supporting evidence in Appendix E with additional statistical testing and alternative popularity metrics.
@@
We also explored whether problem complexity, as measured by the number of expected reasoning steps, correlates with Feedback Friction since prior works [27] show that longer reasoning trajectories may impede problem-solving in LLMs. Additionally, we investigated whether certain questions consistently induced reluctance to feedback across different models because of their inherent characteristics. For both hypotheses, our controlled experiments yielded negative results, with minimal correlation between reasoning complexity and stubbornness, and little overlap in stubborn questions across models. Complete analyses are provided in Appendix F and Appendix G.